A transverse wave on a rope is given by

Y(x,t) = (0.750m)sin2π [(

t

2.50

s) − (

x

0.400

m)]


a) Find the amplitude, period, frequency, wavelength and speed of propagation.

b) Is the wave traveling in the +x or – x direction?


1
Expert's answer
2021-05-25T10:10:23-0400

"y(x,t)=(0.750\ m)\sin 2\pi(\frac{t}{2.5\ s}-\frac{x}{0.400\ m})"


Amplitude "A=0.750\ (m)"


Period "T=2.5\ s"


Frequency "\nu=1/T=1/2.5=0.4\ (Hz)"


Wavelength "\lambda=0.400\ (m)"


Speed of propagation "v=\lambda/v=0.4/0.4=1\ (m/s)"


The wave traveling in the "+x" direction